Which of the following are point-slope equations of the line going through (6, 7) and (2, -1)?
Answer by Marz157(7)

As the name suggests, you need two things to write an equation in point-slope form, a point and the slope.
The slope of a line means how quickly is the line going up or down. You calculate it by seeing how much the Y value changes and divide that by how much the X value changes. You can do this with any two points on the line (Hey, we have two of those!!!)
The difference in Y values here are 7 going to -1, so that's a change of -8. The x values change from 6 to 2, that's a change of -4.
We divide those two values (-8/-4) and get 2. So the slope is 2.
Now, whats the formula for point-slope form? Its this
Y - y = Slope *(X - x)
where the little x and y's are any point on the line. We can use either (6,7) or (2, -1) and it will work.
Y - 7 = 2 * (X - 6)
OR
Y - (-1) = 2 * (X - 2) [Be careful with subtracting the -1 from the Y]
We can double check our answer by plugging in the other point into our equation and make sure that both sides are the same.
Y + 1 = 2 * ( X - 2 ) When X = 6 and Y = 7
7+1=2*(6 - 2)
8 = 2*4
8 = 8